If the rectanggular of the room is twice as long as it is wide. It is diveded into two rectangular part by a partition. If the larger part of the room contains 560 square feet, what is the dimension of the entire room.
Answer by josgarithmetic(39617)

"If the rectanggular of the room", means what? I will use this to mean "a rectangular room".

w and L width and length.
L/w=2, "twice as long as it is wide".
L=2w.
Dimensions of original rectangle are w and 2w.
Partition the room into dimension w and the lengths of each rectangle part is x and y, so that x+y=L. This means, x+y=2w. "Solving" this for y, find y=2w-x. What now is these two rectangles separated by a partition?

One rectangle part is dimensions w by x.
Other rectangle part is dimensions w by 2w-x.
Either wx=560 OR w(2w-x)=560. NOT both.
You seem to have enough information for 1 equation but you have 2 unknown variables.